OpenStudy (anonymous):

integrate the following:

OpenStudy (anonymous):

\[\int\limits_{r1}^{r2} \frac{ dr }{ r^2 }\]

OpenStudy (anonymous):

i have never done a 1/x^2 integral :/

hartnn (hartnn):

don't you know the x^n formula ?

hartnn (hartnn):

1/x^2 = x^ {-2}

hartnn (hartnn):

\(\int x^n dx = x^{n+1}/(n+1)+c\)

hartnn (hartnn):

here n=-2

hartnn (hartnn):

also, x=r :P

OpenStudy (anonymous):

ah ok, i didnt do x^-2 thank you, but how come 1/x is lnIxI?

hartnn (hartnn):

because d/dx of ln x = 1/x so, integral of 1/x = ln x +c

hartnn (hartnn):

that x^n formula is valid for all n, except n=-1

geerky42 (geerky42):

http://www.youtube.com/watch?v=ZGRu55zOOGU

OpenStudy (anonymous):

ah ok :) thank you!

hartnn (hartnn):

welcome ^_^